# GraphCompNet: A Position-Aware Model for Predicting and Compensating Shape Deviations in 3D Printing

### v1

@@ III-A Data Pre-processing
All shape deviation prediction and compensation procedures incorporate a pre-processing step involving metrology and shape registration. The printed parts undergo scanning and quality control using the ATOS scan system and GOM software [35]. The alignment of 3D shapes is attained through a combination of the deep align [36] and iterative closest point (ICP) algorithms [37].
Figure 4: The procedural sequence of the DL prediction engine begins with the conversion of the input CAD into an isometric mesh. Subsequently, graph neural networks are applied to predict shape deviation. The accuracy of the predicted shape deviation is assessed by comparing it with the printed and scanned shapes, utilizing a deformation loss function defined in Equation (9).
Let $\mathbb{C}$ and $\mathbb{S}$ denote the sets of CAD models and scanned point clouds, the training dataset is comprised of pairs of CAD models and corresponding scanned point clouds, denoted as $T={(\mathcal{C}_{1},\mathcal{S}_{1}),\dots,(\mathcal{C}_{n},\mathcal{S}_{n})}$. Here, $\mathcal{C}_{i}\in\mathbb{C}$ represents a CAD mesh, and $\mathcal{S}_{i}\in\mathbb{S}$ represents the scanned point cloud derived from the printed part of $\mathcal{C}_{i}$. Each CAD model $\mathcal{C}_{i}$ is defined as a set of vertices $\mathcal{C}_{i}=\{\mathbf{c}_{1},\dots,\mathbf{c}_{k}\}$, while each scanned point cloud $\mathcal{S}_{i}$ comprises a corresponding set of points $\mathcal{S}_{i}=\{\mathbf{s}_{1},\dots,\mathbf{s}_{k}\}$. The objective of the shape deformation prediction and compensation algorithm is to determine displacement values that predict or compensate for the shape deformation between these two sets of points. The displacement or deformation of a cloud $C_{i}$ is computed at the point level by pairwise matching a point from $C_{i}$ to a point in $S_{i}$, followed by the measurement of the distance between these two points using a specified metric.
@@
$\mathbf{s}_{i}=\mathbf{c}_{i}+g(\mathbf{c}_{i},\mathbf{\theta})\quad\forall\mathbf{c}_{i}\in\mathcal{C}_{i},s_{i}\in\mathcal{S}_{i}$ (3)
@@
In this context, displacement mapping can be directly learned using $g$, and compensation is provided as the inverse function of $g$. However, AM processes typically exhibit non-linear behavior, where critical process variables, such as the temperature field, residual stress and melt pool geometry, spatial correlation within print part and with the neighboring geometries all impact the final outcome. This makes the above model inadequate for proper warpage compensation and resulting in varying shape deformation outcomes [6]. To address this limitation, traditional methods often involve repeating the printing and scanning processes or employing complex model parameterization [9] or B-spline shape registration [6].
@@ III-B GAN-inspired Shape Deviation Compensation
In our approach, the DL prediction engine is trained to recognize shape deviation based on ground truth data (actual printed parts). Subsequently, with the parameters of the DL prediction engine fixed, the DL compensation engine is trained to propose shape compensation plans, which are then evaluated by the DL prediction engine. This iterative process continues as new or updated data becomes available, enhancing the stability of both the DL prediction and compensation engines during training:
$\mathbf{s}_{i}$ $=D(\mathbf{c}_{i};\mathbf{\theta}_{pred})$ (4a)
$\mathbf{c}_{i}^{\prime}$ $=G(\mathbf{c}_{i};\theta_{comp})$ (4b)
$\mathbf{c}_{i}$ $=D(G(\mathbf{c}_{i};\mathbf{\theta}_{comp});\mathbf{\theta}_{pred})$ (4c)
In this context, $G$ denotes the DL compensation engine, responsible for generating compensated CAD vertices. $D$ represents the DL prediction engine, which evaluates shape deviation. The sets of learnable parameters for the DL prediction and compensation engines are denoted by $\mathbf{\theta}_{comp},\mathbf{\theta}_{pred}$, respectively. Specifically, $\mathbf{c}_{i}^{\prime}=G(\mathbf{c}_{i};\mathbf{\theta}_{comp})$ represents the compensated CAD vertex, corresponding to $\mathbf{c}_{i}$, which represents the evaluated shape vertex.
Figure 5: The workflow of the DL compensation engine. The DL compensation engine converts the input CAD into isometric mesh and then applies graph neural networks to find an optimal compensation plan. It passes through a weight-freeze DL prediction engine, then compares the shape deviation results with the input CAD.
@@ III-C1 CAD - Isometric mesh conversion
Geometric primitives (e.g. triangles, rectangles, hexagonal meshes) of the CAD parts are not uniformly shaped but vary in geometry and shape. The precision of the shape deviation prediction/compensation models is proportional to the uniformity and number of geometric primitives. To minimize the effect of the geometric primitives, the CAD parts were converted to isometric meshes. We adopted in-house CAD–mesh conversion software. The isometric mesh $G=\{\mathcal{V},\mathcal{E},\mathcal{F}\}$ can be represented as a graph, where, $\mathcal{V},\mathcal{E},\mathcal{F}$ are sets of vertices, edges, and faces, respectively. Our isometric re-meshing made use of a three-dimensional discrete diffusion approximation. Specifically, input triangular meshes were transformed to a high-density voxelization. The surface voxels, as computed using a 3x3x3 neighborhood, were then progressively wrapped in a manner that approximated a uniform geodesic distance from any other point already sampled. The process started from a random surface voxel and terminated when all of the surface voxels were wrapped. The approximation involved always expanding from a given voxel to immediate neighbors and expanding to diagonal neighbors with a random weight of $1/(4\sqrt[2]{2})$.
@@ III-C2 Graph neural network
Graph neural network is specifically designed to apply convolution to non-Euclidean data such as CADs, with nodes, edges, and faces form an isometric mesh. There are several graph convolution techniques to define $h_{\theta}$. In this work, we used edge convolution-based graph neural networks [31]. Given an isometric mesh graph $G=\left(\mathcal{V},\mathcal{E}\right)$,
@@
where, (||)($\cdot$|$\cdot$)( ⋅ | | ⋅ ) is a concatenation operation between two tensors, $x_{j}-x_{i}$ captures the local neighborhood information, i.e. the graph neural network provides not only global shape but also local shape information.
Using isometric mesh as a graph provides two advantages. First of all, it is computationally cheaper than using $k$-nn graphs as it is pre-determined by existing mesh structures. Second, geometric primitives, that forms an isometric mesh, naturally encode the parts’ surface geometry, such that it gives a better representation to predict shape deviation. With the standard activation layer (e.g. ReLU) and the edge convolution $h_{\theta}$ defined in Equation (5), edge convolution layers are defined as:
@@
The detailed implementation of graph neural networks in the DL compensation/prediction engines will be explained in the result section.
III-C3 Evaluation with ground truth
The predictions from the graph neural network module must be evaluated by a proper metric. Without a valid metric, the DL prediction engine cannot be trained to achieve high geometric precision and accuracy. The standard $L_{2}$ loss (a.k.a mean square loss) is used to compute regression between the isometric mesh and the scanned point clouds. Let $C_{1},S_{1}$ be the sets of vertices and corresponding point clouds from the CAD and scanned part, respectively. The standard $L_{2}$ loss is defined as follows
$L2\_\text{Loss}(\mathcal{C}_{1},\mathcal{S}_{1})\ =\ \frac{1}{n}\sum_{i=1}^{n}{(\mathbf{c}_{i}-\mathbf{s}_{i})^{2}}$ (7)
where, $\mathbf{c}_{i}\in\mathcal{C}_{1}$ and $\mathbf{s}_{j}\in\mathcal{S}_{1}$.
However, $L_{2}$ loss does not provide shape consistency, such that it may cause some oscillations or other irregular patterns. In order to preserve shape consistency, we introduce the Chamfer loss function. Chamfer loss is defined as follows
[…]
$\sum_{\mathbf{c}_{i}\in\mathcal{C}_{1}}\min_{\mathbf{s}_{i}\in\mathcal{S}_{1}}\|\mathbf{c}_{i}-\mathbf{s}_{i}\|+\sum_{\mathbf{s}_{i}\in\mathcal{S}_{1}}\min_{\mathbf{c}_{i}\in\mathcal{C}_{1}}\|\mathbf{c}_{i}-\mathbf{s}_{i}\|$ (8)
Chamfer distance computes the sum of the smallest distances between each element in $\mathcal{C}_{1}$ and $\mathcal{S}_{1}$, thus penalizing shape inconsistency between $\mathcal{C}_{1}$ and $\mathcal{S}_{1}$. In summary, our loss function is defined as:
[…]
@@ III-D DL compensation Engine
Figure 6: Performance of the trained prediction engine on the sample test bar data. The heatmap (units in mm) displays the geometric distance between two 3D difference pairs: the top heatmap shows the difference between the original CAD design and the scanned actual print geometry, while the bottom heatmap shows the difference between the trained prediction engine and the scanned print geometry.The trained prediction engine’s predicted deviation closely aligns with the ground truth.
@@
$L2\_\text{Loss}(\mathcal{C}_{1},\mathcal{O}_{1})\ =\ \frac{1}{n}\sum_{i=1}^{n}{(\mathbf{c}_{i}-\mathbf{o}_{i})^{2}}$ (10)
@@
$\sum_{\mathbf{c}_{i}\in\mathcal{C}_{1}}\min_{\mathbf{o}_{i}\in\mathcal{O}_{1}}\|\mathbf{c}_{i}-\mathbf{o}_{i}\|+\sum_{\mathbf{o}_{i}\in\mathcal{O}_{1}}\min_{\mathbf{c}_{i}\in\mathcal{C}_{1}}\|\mathbf{c}_{i}-\mathbf{o}_{i}\|$ (11)
[…]
where, $\mathbf{o}_{i}\in\mathcal{O}_{1}$, $\mathbf{c}_{j}\in\mathcal{C}_{1}$ and:
$\mathbf{c}_{i}^{\prime}$ $=G(\mathbf{c}_{i};\theta_{comp})$ (13a)
$\mathbf{o}_{i}$ $=D(G(\mathbf{c}_{i};\mathbf{\theta}_{comp});\mathbf{\theta}_{pred})$ (13b)
@@ IV Case Studies
As discussed in Section 2, existing approaches show limited efficiency in handling complex topologies and face challenges in closed-loop applications for batch production, primarily due to the high cost of repetitive experimental validation and the failure to account for position-dependent variables. In this context, we highlight the effectiveness of the proposed Predictor-Compensator framework.
As illustrated in Figure 1 and Figure 9 (a), large parts often experience significant warping due to the varying thermal conditions across different regions of the build chamber. To assess the impact of nesting (the strategic arrangement of objects within a 3D printer’s build chamber) on part warpage, in Case Study A, we validated the approach by evaluating a nested bucket containing 140 bar-shaped parts positioned at various positions within the print chamber. In Case Study B, we applied the proposed computational framework to a more complex geometry, referred to as Molded Fiber data, highlighting the proposed method’s versatility across different nesting orientations and angles. Scanning and 3D reconstruction of the printed designs were conducted using the GOM Suite system (GOM, Germany). Both the DL prediction and compensation engines employ identical graph neural network architectures, comprising four edge convolution layers (see Figure 4 and 5). The proposed architecture was implemented using the PyTorch and PyTorch Geometry libraries. Training of the networks employed the Adam optimizer with a learning rate of $0.001$ and 1000 epochs. All experiments were conducted using a single GeForce RTX 3090 GPU with 24GB VRAM memory.

### v3

@@ III-A Data Pre-processing
All shape deviation prediction and compensation procedures incorporate a pre-processing step involving metrology and shape registration. The printed parts undergo scanning and quality control using the ATOS scan system and GOM software [VAGOVSKY20151198]. To enable point-level comparison with the nominal CAD geometry, we align the scanned mesh and CAD model using a combination of Deep Align [JIMENEZMORENO2021107712] and iterative closest point (ICP) algorithms [article_icp]. After registration, both the CAD and scanned meshes are uniformly resampled to the same number of points, $k$, to standardize the input resolution for model training. This resampling ensures consistency in shape representation, though it does not assume a strict one-to-one correspondence between CAD and scan vertices.
Let $\mathbb{C}$ and $\mathbb{S}$ denote the sets of CAD models and scanned point clouds, the training dataset is comprised of pairs of CAD models and corresponding scanned point clouds, denoted as $T={(\mathcal{C}_{1},\mathcal{S}_{1}),\dots,(\mathcal{C}_{n},\mathcal{S}_{n})}$. Here, $\mathcal{C}_{i}\in\mathbb{C}$ represents a CAD mesh, and $\mathcal{S}_{i}\in\mathbb{S}$ represents the scanned point cloud derived from the printed part of $\mathcal{C}_{i}$. Each CAD model $\mathcal{C}_{i}$ is defined as a set of vertices $\mathcal{C}_{i}=\{\mathbf{c}_{1},\dots,\mathbf{c}_{k}\}$, while each scanned point cloud $\mathcal{S}_{i}$ comprises a corresponding set of points $\mathcal{S}_{i}=\{\mathbf{s}_{1},\dots,\mathbf{s}_{k}\}$. The objective of the shape deformation prediction and compensation algorithm is to determine displacement values that predict or compensate for the shape deformation between these two sets of points. The displacement or deformation of a cloud $C_{i}$ is computed at the point level by pairwise matching a point from $C_{i}$ to a point in $S_{i}$, followed by the measurement of the distance between these two points using a specified metric. The simulated or predicted point cloud of the printed part with deviations is annotated as $\hat{\mathcal{S}}_{i}=\{\hat{\mathbf{s}}_{1},\dots,\hat{\mathbf{s}}_{k}\}$.
@@
$\hat{\mathbf{s}}_{i}=\mathbf{c}_{i}+g(\mathbf{c}_{i},\mathbf{\theta})\quad\forall\mathbf{c}_{i}\in\mathcal{C}_{i},\hat{\mathbf{s}}_{i}\in\hat{\mathcal{S}}_{i}$ (3)
@@
In this context, displacement mapping can be directly learned using $g$, and compensation is provided as the inverse function of $g$. However, AM processes typically exhibit non-linear behavior, where critical process variables, such as the temperature field, residual stress and melt pool geometry, spatial correlation within print part and with the neighboring geometries all impact the final outcome. This makes the above model inadequate for proper warpage compensation and resulting in varying shape deformation outcomes [schmutzler2016compensating]. To address this limitation, traditional methods often involve repeating the printing and scanning processes or employing complex model parameterization [ferreira2019automated] or B-spline shape registration [schmutzler2016compensating].
@@ III-B GAN-inspired Shape Deviation Compensation
In our approach, the DL prediction engine is trained to recognize shape deviation based on ground truth data (actual printed parts). Subsequently, with the parameters of the DL prediction engine fixed, the DL compensation engine is trained to propose shape compensation plans, which are then evaluated by the DL prediction engine. This iterative process continues as new or updated data becomes available, enhancing the stability of both the DL prediction and compensation engines during training. Below Eq. 4a indicates the predicted shape deviation for original CAD vertex, Eq. 4b indicates the compensated CAD vertex from the compensation engine and Eq. 4c indicates the predicted shape deviation for the compensated vertex.
$\hat{\mathbf{s}}_{i}$ $=D(\mathbf{c}_{i};\mathbf{\theta}_{pred})$ (4a)
$\mathbf{c}_{i}^{\prime}$ $=G(\mathbf{c}_{i};\mathbf{\theta}_{comp})$ (4b)
$\hat{\mathbf{s}}_{i}^{\prime}$ $=D(\mathbf{c}_{i}^{\prime};\mathbf{\theta}_{pred})$ (4c)
Notation Explanation:
  • $\mathbf{c}_{i}$: The original CAD vertex before compensation.
  • $\mathbf{c}_{i}^{\prime}$: The compensated CAD vertex proposed by the compensation engine $G$.
  • $\hat{\mathbf{s}}_{i}$: The predicted position of the printed part surface corresponding to the original CAD vertex $\mathbf{c}_{i}$ as estimated by the prediction model $D$. This does not represent a difference vector only, but rather the simulated location of the deformed surface point.
  • $\hat{\mathbf{s}}_{i}^{\prime}$: The predicted position of the printed part surface corresponding to the compensated CAD vertex $\mathbf{c}_{i}^{\prime}$, also estimated by the prediction model $D$. This reflects the expected outcome after compensation and includes modeled deformation effects.
  • $D(\cdot;\mathbf{\theta}_{pred})$: The DL prediction engine, parameterized by $\mathbf{\theta}_{pred}$.
  • $G(\cdot;\mathbf{\theta}_{comp})$: The DL compensation engine, parameterized by $\mathbf{\theta}_{comp}$.
The learning process aims to optimize the parameters $\mathbf{\theta}_{comp}$ of the DL compensation engine $G$ to minimize the shape deviation between the compensated vertices $\mathbf{c}_{i}^{\prime}$ and the actual scanned points $\mathbf{s}_{i}$ by minimizing the deformation loss.
Figure 5: The workflow of the DL compensation engine. The DL compensation engine converts the input CAD into isometric mesh ($\mathcal{C}_{i}$) and then applies graph neural networks to find an optimal compensation plan. It passes through a weight-freeze DL prediction engine, then compares the shape deviation results with the input CAD.
@@ III-C1 CAD - Isometric mesh conversion
Geometric primitives in CAD models—such as triangles, rectangles, or hexagonal meshes—are inherently irregular, exhibiting significant variability in shape and size. While this non-uniformity offers a compact geometric description, it can negatively impact the accuracy of shape deviation prediction and compensation models. Specifically, the precision of these models is strongly correlated with the uniformity and density of the underlying geometric primitives. To mitigate the influence of irregular primitive shapes, we convert the original CAD parts into isometric meshes using our in-house CAD-to-mesh conversion software. The resulting isometric mesh $G=\{\mathcal{V},\mathcal{E},\mathcal{F}\}$ is represented as a graph, where, $\mathcal{V},\mathcal{E},\mathcal{F}$ are sets of vertices, edges, and faces, respectively. Our re-meshing procedure leverages a three-dimensional discrete diffusion approximation. The input triangular mesh is first transformed into a high-density voxel grid. Surface voxels are identified by analyzing their 3×3×3 neighborhoods. Starting from a randomly selected surface voxel, the mesh is progressively “wrapped” by expanding to adjacent voxels, including diagonal neighbors weighted probabilistically by $1/(4\sqrt[2]{2})$. This expansion continues until all surface voxels are encompassed, resulting in a mesh with near-uniform vertex spacing and geodesic distances.
Using isometric meshes offers several advantages compared to working with the original non-uniform meshes. The uniform vertex distribution reduces bias introduced by irregular primitive sizes and shapes, facilitating more consistent and stable graph-based learning. This uniformity leads to more stable graph-based learning and better predictive accuracy for geometric deviation and compensation. Without re-meshing, heterogeneous local geometries can hinder model performance and generalization.
@@ III-C2 Graph neural network
Graph neural networks (GNNs) enable convolution over non-Euclidean data such as CAD surfaces represented by isometric meshes. In this work, we employ edge convolution-based GNNs [wang2019dynamic], which operate on a mesh graph $G=\left(\mathcal{V},\mathcal{E}\right)$ as follows:
@@
where, $(\cdot||\cdot)$ is a concatenation operation between two tensors, $x_{j}-x_{i}$ captures the local neighborhood information, i.e. the graph neural network provides not only global shape but also local shape information.
Using isometric mesh as a graph provides two advantages. First of all, it is computationally cheaper than using $k$-$n$ graphs as it is pre-determined by existing mesh structures. Second, geometric primitives, that forms an isometric mesh, naturally encode the parts’ surface geometry, such that it gives a better representation to predict shape deviation. With the standard activation layer (e.g. ReLU) and the edge convolution $h_{\theta}$ defined in Equation (5), edge convolution layers are defined as:
@@
The detailed implementation of graph neural networks in the DL compensation / prediction engines are explained in the following sections.
III-C3 Training the DL Prediction Engine
The DL prediction engine $D(\mathbf{c}_{i};\mathbf{\theta}_{pred})$ is trained by minimizing the deformation loss between the predicted shape deviation $\hat{\mathbf{s}}_{i}$ and the actual scanned point $\mathbf{s}_{i}$. The optimization problem is defined as:
$\min_{\mathbf{\theta}_{pred}}\quad\text{Deformation\_loss}$
$\text{Deformation\_loss}=L2\_\text{Loss}+\text{Chamfer\_loss}$ (7)
The standard $L_{2}$ loss (a.k.a mean square loss) is used to compute regression between the isometric mesh and the deformation, represented by the scanned point clouds. Let $\mathcal{S},\hat{\mathcal{S}}$ be the sets of vertices and corresponding point clouds from scanned part and the model predicted deviation, respectively. The standard $L_{2}$ loss is defined as follows:
$L2\_\text{Loss}\ =\frac{1}{n}\sum_{i=1}^{n}\|\hat{\mathbf{s}}_{i}-\mathbf{s}_{i}\|^{2}$ (8)
However, $L_{2}$ loss does not provide shape consistency, such that it may cause some oscillations or other irregular patterns. In order to preserve shape consistency, we introduce the Chamfer loss function. Chamfer distance computes the sum of the smallest distances between each element in $\hat{\mathcal{S}}$ and $\mathcal{S}$, thus penalizing shape inconsistency between $\hat{\mathcal{S}}$ and $\mathcal{S}$. It is defined as:
$\text{Chamfer\_loss}=\sum_{\hat{\mathbf{s}}_{i}\in\hat{\mathcal{S}}}\min_{\mathbf{s}_{j}\in\mathcal{S}}\|\hat{\mathbf{s}}_{i}-\mathbf{s}_{j}\|+\sum_{\mathbf{s}_{j}\in\mathcal{S}}\min_{\hat{\mathbf{s}}_{i}\in\hat{\mathcal{S}}}\|\mathbf{s}_{j}-\hat{\mathbf{s}}_{i}\|$ (9)
Here:
  • $\mathbf{c}_{i}$: Original CAD vertex.
  • $\mathbf{s}_{i}$: Actual scanned point (ground truth).
  • $\hat{\mathbf{s}}_{i}=D(\mathbf{c}_{i};\mathbf{\theta}_{pred})$: The predicted position of the printed part surface corresponding to the original CAD vertex $\mathbf{c}_{i}$, as estimated by the prediction model $D$. This includes anticipated shape deviations due to process-induced distortion.
  • $\mathbf{\theta}_{pred}$: Parameters of the DL prediction engine to be optimized during training.
@@ III-D DL compensation Engine
Figure 6: Prediction performance on the sample test bar, visualized from a top-down (planar) view. The heatmaps (units in $m$) represent geometric deviation magnitudes across the build plate. The top heatmap shows the deviation between the original CAD model and the scanned printed part, capturing actual deformation. The bottom heatmap compares the trained prediction engine’s output to the scanned print, illustrating the model’s accuracy. The non-uniform distortion observed in the physical print is well approximated by the prediction engine, demonstrating its ability to capture position-dependent deformation behavior.
@@
$\text{Chamfer\_loss}(\mathcal{C}_{1},\hat{\mathcal{S}}_{1}^{\prime})=$
$\sum_{\mathbf{c}_{i}\in\mathcal{C}}\min_{\hat{\mathbf{s}}_{j}^{\prime}\in\hat{\mathcal{S}}^{\prime}}\|\mathbf{c}_{i}-\hat{\mathbf{s}}_{j}^{\prime}\|+\sum_{\hat{\mathbf{s}}_{j}^{\prime}\in\hat{\mathcal{S}}^{\prime}}\min_{\mathbf{c}_{i}\in\mathcal{C}}\|\hat{\mathbf{s}}_{j}^{\prime}-\mathbf{c}_{i}\|$ (11)
$\text{Deformation loss}=L2\_\text{Loss}+\text{Chamfer\_loss}$
Here:
  • $\mathbf{\theta}_{comp}$: Parameters of the DL compensation engine to be optimized during training.
  • $\mathbf{\theta}_{pred}$: Parameters of the DL prediction engine, which is fixed during the DL compensation engine training.
@@ IV Case Studies
As noted in Section II, prior approaches struggle with complex topologies and are limited in closed-loop batch production due to high experimental costs and lack of position-specific modeling. To demonstrate the effectiveness of our Predictor-Compensator framework, we conducted case studies using real parts printed on an HP Multi Jet Fusion (MJF) system with PA12, a thermoplastic known for strength and dimensional stability. This setup reflects a common industrial environment, underscoring the framework’s practical relevance.
As shown in Figures 1 and 8(a), large parts often exhibit warping from thermal gradients across the build chamber. To evaluate the effect of nesting (the strategic arrangement of objects within a 3D printer’s build chamber) on distortion, Case Study A involved a nested bucket containing 140 bar-shaped parts printed at different positions. Case Study B extended the framework to a complex geometry (Molded Fiber), demonstrating adaptability to varying part orientations and shapes. Printed parts were scanned and reconstructed using the GOM Suite (GOM, Germany). Both the deviation prediction and compensation modules use the same graph neural network architecture with four edge convolution layers (see Figures 4 and 5), implemented in PyTorch and PyTorch Geometric. Models were trained using the Adam optimizer (learning rate = 0.001) for 1000 epochs on a single NVIDIA RTX 3090 GPU (24GB VRAM).
Figure 7: Performance of the proposed deep learning–based compensation engine on test data. The heatmaps (units in $m$) visualize geometric deviation magnitudes between different 3D shape comparisons. Specifically, the plots include deviations between the nominal CAD model and the scanned printed part, the nominal CAD and the compensated CAD geometry, the nominal CAD and the predicted position of the print part surface for the model proposed compensated CAD geometry. These visualizations illustrate how well the compensation strategy reduces geometric errors relative to the original design.